# Fermi Machine — Quantum Many-Body Solver Derived from Correspondence between Noninteracting and Strongly Correlated Fermions

### v1

@@ 4 Fermi Machine as a Quantum Neural Network
$\tilde{d}^{(m)\dagger}_{\ell,\sigma}$$\rightarrow$ $\sum_{\ell^{\prime},\sigma^{\prime}}\Theta_{\sigma^{\prime},\sigma}^{(m)}(\ell^{\prime}-\ell)d^{(m-1)\dagger}_{\ell,\sigma}(-)^{n_{\tilde{d}^{(m-1)},\ell^{\prime},\sigma^{\prime}}},$ (88)
$\rightarrow$ $\Xi^{(m)}_{\sigma}(q,\ell)c_{\ell,\sigma}^{\dagger},$
$\Xi^{(m)}_{\sigma}(q,\ell)$ $=$ $\prod_{n=1}^{m}[\sum_{\ell_{n},\sigma_{n}}\Theta_{\sigma_{n},\sigma}^{(m-n+1)}(q,\ell_{n}-\ell)(-)^{mn_{\ell_{n},\sigma_{n}}}],$
$\Xi^{(0)}_{\sigma}(\ell)$ $=$ $1,$ (90)
@@
$|\Psi_{\sigma}^{\rm Hub}(q)\rangle$ $=$ $\sum_{m=0}^{M}\alpha_{m,q}\sum_{\ell,\ell^{\prime},\sigma^{\prime}}\frac{e^{iq\ell}}{\sqrt{N_{s}}}\Xi_{\sigma^{\prime},\sigma}^{(m)}(q,\ell^{\prime}-\ell)c_{\ell,\sigma}^{\dagger}|0\rangle.$
Figure 2: Illustration of the architecture of Fermi machine network extended from Fig. 1 by introducing momentum dependent (spatially extended) hybridization.
@@
  • For given parameters in Eqs. (55)-(58), determine $\alpha_{m,k}$ for the ground state of the MCFM and then $|\Psi_{0}^{\rm Hub}\rangle$ in the form of Eqs. (67) and (69) with Eqs. (70), (LABEL:Psi_Hub_qsigma2), (86) or (LABEL:Psi_Hub_qsigma5).
@@
  • Optimize the variational parameters in Eqs. (LABEL:eq:MCFM_Hc)-(58) to lower $E_{0}$ until convergence. For the optimization, the natural gradient [37] (stochastic reconfiguration [38]) can be used.
@@
Further refinements discussed above may also be applied as well. As a simple starting point one may impose a constraint of the “quenched randomness” for the construction of the MCFM, which means that the MCFM paramters are taken not to depend on the layer for the same site or bond in the corresponding Hubbard model.
@@ 5 Benchmark for 4-site Hubbard model
Table 2: Comparison of estimated ground-state energies per site $E_{0}$ of the hole doped 4-site Hubbard model between the exact diagonalization and Fermi machine with list of an example of optimized variational parameters for $U=8,t=1$ with 1 up and down doped hole each. Parameters not listed in the table are $M=1$, $M^{\prime}=1$, $t_{\tilde{c}}=t_{\tilde{d}}=1$, $t_{\tilde{d},i,j,\sigma}^{(M+1)}=0$, and $\mu_{\tilde{d},i}^{(M+1)}=\mu_{\tilde{d},i}^{(M+1)}=8.0$. The nonlocal hybridization is taken in a simple form as $\Lambda^{(M)}_{\sigma,\sigma^{\prime}}[k]=\lambda^{(M)}_{\sigma,\sigma^{\prime}}\cos[k]$ so that the interlayer hybridization is extended simply to the nearest neighbor pair of $\{i,j\}$ in Eq. (84). Accordngly, we take $\Theta_{\sigma,\sigma^{\prime}}(i-j)=1$ only for the nearest neighbor pair of ${i,j}$ for any combination of $\sigma$ and $\sigma^{\prime}$ and zero otherwise in Eq. (LABEL:Psi_Hub_qsigma2).
@@ 5.2 Doped Case
Table 2 shows the doped case (1 up and 1 down holes doped) for $U=8$ and $t=1$. Here, the nonlocal hybridization term is necessary to reproduce the exact result. Still small numbers of parameters are enough to reproduce the exact result.
@@ 6 Discussion, Summary and Outlook
The present correspondence relation establishes the relation between the “bulk” noninteracting multi-component fermions and the “edge” strongly correlated fermions. This has a conceptual similarity with the holographic correspondence in the AdS-CFT [41, 42, 43] and the bulk-edge correspondence in topological materials [44, 45]. Further deep connections to these two frontiers are an interesting future issue.
@@ Appendix A Repsentability of the ground state
This can be interpreted as the representation by a linear combination of $M$ Slater determinants, where the $m$-th one is $|\Psi_{\sigma,m}^{\rm Hub}\rangle$, aside from the entanglement of $\sigma$ and $-\sigma$ electrons contained in the odd $m$ terms. This exchange becomes justified in the following setup: By taking $\alpha_{m,q}=0$ for the odd $m$, which is the case of $\mu_{\tilde{d},i}^{(m)}=\infty$ for odd $m$, the even $m$ layers are indeed decoupled each other and each multi-particle state becomes a simple Slater determinant consisting of the single layer component with even $m$. Then if the Hamiltonian parameters of each layer is taken in such a way that the ground-state energy of each single-layer Slater determinant is degenerate with those of different layers, we may take any linear combination of them, generating a multi-Slater determinant. (Note that in this case, the weights of the linear combination are the variational parameters as well.) The construction of the MCFM Hamiltonian parameters to satisfy the degeneracy is of course possible, because the MCFM energy and occupied momenta at each layer can be freely tuned by taking appropriate dispersion at each layer. Since the complete set of the Hilbert space can be expanded by linear combinations of fermion determinants, this is a formal proof of the representability of the Hubbard model (more generally, any interacting lattice fermion system).
Of course, the number of Slater determinants for the complete representability increases rapidly with increasing system size if one follows this naive muli-determinant representation and this proof is just formal, while practically, the increase can be suppressed by a proper choice of the odd $m$ contribution as well as by the spatially off-diagonal hybridization, as is dicussed from Eq.(83) through Eq. (LABEL:Psi_Hub_qsigma5).

### v2

@@ 4 Fermi Machine as a Quantum Neural Network
$\frac{e^{iq\ell}}{\sqrt{N_{s}}}\tilde{d}^{(m)\dagger}_{\ell,\sigma}$$\rightarrow$ $\sum_{\ell^{\prime},\sigma^{\prime}}\frac{e^{iq\ell^{\prime}}}{\sqrt{N_{s}}}\Theta_{\sigma^{\prime},\sigma}^{(m)}(q,\ell^{\prime}-\ell)d^{(m-1)\dagger}_{\ell,\sigma}$
×$\times(-)^{n_{\tilde{d}^{(m-1)},\ell^{\prime},\sigma^{\prime}}}$ (87)
$\rightarrow$ $\Xi^{(m)}_{\sigma}(q,\ell)c_{\ell,\sigma}^{\dagger},$ (88)
$\Xi^{(m)}_{\sigma}(q,\ell)$ $=$ $\prod_{n=1}^{m}\large[\sum_{\ell_{n},\sigma_{n}}\frac{e^{iq\ell_{n}}}{\sqrt{N_{s}}}\Theta_{\sigma_{n},\sigma}^{(m-n+1)}(q,\ell_{n}-\ell)$ (89)
×$\times(-)^{n_{\ell_{n},\sigma_{n}}}\large],$
$\Xi^{(0)}_{\sigma}(\ell)$ $=$ $\frac{e^{iq\ell}}{\sqrt{N_{s}}},$ (90)
@@
$|\Psi_{\sigma}^{\rm Hub}(q)\rangle$ $=$ $\sum_{m=0}^{M}\alpha_{m,q}\sum_{\ell}\Xi_{\sigma}^{(m)}(q,\ell)c_{\ell,\sigma}^{\dagger}|0\rangle.$ (91)
Figure 2: (Color online): Illustration of the architecture of Fermi machine network extended from Fig. 1 by introducing momentum dependent (spatially extended) hybridization.
@@
  • For given parameters in Eqs. (55)-(58), determine $\alpha_{m,k}$ for the ground state of the MCFM and then $|\Psi_{0}^{\rm Hub}\rangle$ in the form of Eqs. (67) and (69) with Eqs. (70), (85), (86) or (91).
@@
  • Optimize the variational parameters in Eqs. (LABEL:eq:MCFM_Hc)-(58) to lower $E_{0}$ until convergence. For the optimization, the natural gradient [37, 38] (stochastic reconfiguration [39]) can be used.
@@
Further refinements discussed above may also be applied as well. As a simple starting point one may impose a constraint of the “quenched randomness” for the construction of the MCFM, which means that the MCFM paramters are taken not to depend on the layer index for the same site or bond.
@@ 5 Benchmark for 4-site Hubbard model
Table 2: Comparison of estimated ground-state energies per site $E_{0}$ of the hole doped 4-site Hubbard model between the exact diagonalization and Fermi machine with list of an example of optimized variational parameters for $U=8,t=1$ with 1 up and down doped hole each. Parameters not listed in the table are $M=1$, $M^{\prime}=1$, $t_{\tilde{c}}=t_{\tilde{d}}=1$, $t_{\tilde{d},i,j,\sigma}^{(M+1)}=0$, and $\mu_{\tilde{d},i}^{(M+1)}=\mu_{\tilde{d},i}^{(M+1)}=8.0$. The nonlocal hybridization is taken in a simple form as $\Lambda^{(M)}_{\sigma,\sigma^{\prime}}[k]=\lambda^{(M)}_{\sigma,\sigma^{\prime}}\cos[k]$ so that the interlayer hybridization is extended simply to the nearest neighbor pair of $\{i,j\}$ in Eq. (84). Accordngly, we take $\Theta_{\sigma,\sigma^{\prime}}(i-j)=1$ only for the nearest neighbor pair of ${i,j}$ for any combination of $\sigma$ and $\sigma^{\prime}$ and zero otherwise in Eq. (85).
@@ 5.2 Doped Case
Table 2 shows the doped case (1 up and 1 down holes doped) for $U=8$ and $t=1$. Here, the nonlocal hybridization term is necessary to reproduce the exact result. Still small numbers of parameters are enough.
@@ 6 Discussion, Summary and Outlook
The present correspondence establishes the relation between the “bulk” noninteracting multi-component fermions and the “edge” strongly correlated fermions. This has a conceptual similarity in other correspondence in physics such as the holographic correspondence in the AdS-CFT [42, 43, 44], the bulk-edge correspondence in topological materials [45, 46] and the mapping of $d$-dimentional quantum systems to $d+1$-dimensional classical ones in the path integral, though physical contents are substantially different. Further deep connections to these three frontiers are interesting future issues.
@@ Appendix A Repsentability of the ground state
This can be interpreted as the representation by a linear combination of $M$ Slater determinants, where the $m$-th one is $|\Psi_{\sigma,m}^{\rm Hub}\rangle$, aside from the entanglement of $\sigma$ and $-\sigma$ electrons contained in the odd $m$ terms. This exchange becomes justified in the following setup: By taking $\alpha_{m,q}=0$ for the odd $m$, which is the case of $\mu_{\tilde{d},i}^{(m)}=\infty$ for odd $m$, the even $m$ layers are indeed decoupled each other and each multi-particle state becomes a simple Slater determinant consisting of the single-layer component with even $m$. Then if the Hamiltonian parameters of each even $m$th layer is taken in such a way that the ground-state energy of each single-layer Slater determinant is degenerate with those of different even $m$th layers, we may take any linear combination of them, generating a multi-Slater determinant. (Note that in this case, the weights of the linear combination are the variational parameters as well.) The construction of the MCFM Hamiltonian parameters to satisfy the degeneracy is of course possible, because the MCFM energy and occupied momenta at each layer can be freely tuned by taking appropriate dispersion at each layer. Since the complete set of the Hilbert space can be expanded by linear combinations of fermion determinants, this is a formal proof of the representability of the Hubbard model (more generally, any interacting lattice fermion system).
Of course, the number of Slater determinants for the complete representability increases rapidly with increasing system size if one follows this naive muli-determinant representation and this proof is just formal, while practically, the increase can be suppressed by a proper choice of the odd $m$ contribution as well as by the spatially off-diagonal hybridization, as is discussed from Eq.(83) through Eq. (91).